OpenStudy (anonymous):

A cereal factory has three machines for dispensing cereal into bags. A test is done in which 8 bags are filled by each machine and the bags are weighed. The results in grams: Machine 1: 309, 311.5, 313, 310, 308, 314, 312, 310.5 Machine 2: 310, 309, 314, 311, 309, 314.5, 312.5, 310 Machine 3: 309, 310.5, 313.5, 315.5, 310, 311, 312, 309.5 Which statement is true about the mean output of the three machines?Answer: a. Machine 1 has the highest b. Machine 2 has the highest c. Machine 3 has the highest d. All three machines have the same mean e. Two machines sh

OpenStudy (lukebluefive):

Take the average for each machine: Machine 1 Average = 311 Machine 2 Average = 311.25 Machine 3 Average = 311.75

OpenStudy (lukebluefive):

Average = (Sum of all values) / (Number of values) Number of values = 8 in this case

OpenStudy (anonymous):

umm you have to explain better then that cause im not that smart i dropped out in 7th grade bro i didnt learn none of that

OpenStudy (anonymous):

workin on my ged

OpenStudy (lukebluefive):

Ah, well, hope that goes well. I've heard it's tough. Sorry, let me explain better this time. First, sum all of the values for each machine, individually: Machine 1 Total Dispensed = 2488 grams Machine 2 Total Dispensed = 2490 grams Machine 3 Total Dispensed = 2491 grams

OpenStudy (lukebluefive):

To get the average, or mean, you divide these numbers by the total number of values, or bags, in this case, which number 8: Machine 1 Average Dispensed = 2488 / 8 = 311 Machine 2 Average Dispensed = 2490 / 8 = 311.25 Machine 3 Average Dispensed = 2491 / 8 = 311.75

OpenStudy (lukebluefive):

The question is asking which machine outputs the most cereal on average.

OpenStudy (lukebluefive):

Does this make sense?

OpenStudy (anonymous):

are you good in language arts

OpenStudy (anonymous):

Pablo needs to buy speaker wire for his stereo. He wants the two speakers to be 28 feet from the stereo's central console. If the speaker wire is $3.50 per foot, what is the minimum that Pablo will have to spend on speaker wire? Answer: a. $31.50 b. $49.00 c. $56.00 d. $98.00 e. $196.00

OpenStudy (lukebluefive):

I'm above average, though it's not my specialty. What do you need help with?

OpenStudy (anonymous):

the question i just sent forget the language arts for right now

OpenStudy (lukebluefive):

Well, since there are two speakers, each 28 feet from the central console, that'd be: 2 x 28 feet x $3.50 = 28 feet x $7 = ? Assuming that two speakers can't share the same wire, which seems like a realistic assumption to me.